OpenStudy (anonymous):

If log (a) x = 7, what is the value of log a (a/ x^2) ? Thank you!

OpenStudy (anonymous):

It can be simplified to log(a) a - 2 log (a)x= 1-14= -13

OpenStudy (campbell_st):

well apply the log law for division and you get \[\log_{a}(\frac{a^1}{x^2}) = \log_{a} (a^1)- \log_{a}(x^2)\] and you should know \[\log_{a}(a) = 1\] or is equal to the power so its \[1 - \log_{a}(x^2)\] now apply the log law for indexes to the 2nd term \[\log(m^b) = b \times \log(m)\] then you should have an expression in terms of log(x) hope it helps

OpenStudy (anonymous):

Thank you! What do m and b represent?

OpenStudy (campbell_st):

in you're question its x^2

OpenStudy (anonymous):

So m represents x and b represents root 2?

OpenStudy (anonymous):

I'm still not quite sure how to solve this.

OpenStudy (campbell_st):

well if you look at it the simplification using log laws becomes \[\log_{a}(a^1) - \log_{a}(x^2) = 1 - 2 \times \log_{a}(x)\] and you know the value of log(x) so substitute it and then find the value.

OpenStudy (anonymous):

I just tried to solve it and got 1.535, but I don't think thats right.

OpenStudy (campbell_st):

ok... so you know \[\log_{a}(x) = 7\] that piece of information is given in the question so then \[1 - 2 \times \log_{a}(x) = 1 - 2 \times 7 \]

OpenStudy (anonymous):

Oh! So it's -13?

OpenStudy (campbell_st):

thats it